2005 TEX Users Group Election

Election Notice

The terms of the TUG President and of eight members of the Board of Directors will be open as of the TUG 2005 Annual Meeting, which will be held in August 2005, in Wuhan, China.

The current President, Karl Berry, has stated that he will stand for re-election. The directors whose terms will expire in 2005 are: Steve Grathwohl, Jim Hefferon, Arthur Ogawa, Gerree Pecht, Steve Peter, and Mike Sofka. Continuing directors, with terms ending in 2007: Barbara Beeton, Kaja Christiansen, Susan DeMeritt, Ross Moore, Cheryl Ponchin, Samuel Rhoads, and Philip Taylor.

The election to choose the new President and Board members will be held in the spring of 2005. Nominations for these openings are now being invited.

The TUG Bylaws provide that “Any member may be nominated for election to the office of TUG President/to the Board by submitting a nomination petition in accordance with the TUG Election Procedures. Election … shall be by written mail ballot of the entire membership, carried out in accordance with those same Procedures.”

The term of a member of the Board is four years; President is two years.

The name of any member may be placed in nomination for election to one of the open offices by submission of a petition, signed by two other members in good standing, to the TUG office. A nomination form follows this announcement; forms may also be obtained from the TUG office, and electronically

via the TUG web site at http://tug.org/election.

Along with a nomination form, each candidate is asked to supply a a short biography, a statement of intent, and (optional) passport-size photograph to be included with the ballot. The biography and statement of intent together may not exceed 400 words. Also, a candidate’s membership dues for 2005 are expected to be paid by the nomination deadline.

The deadline for receipt at the TUG office of nomination forms and ballot information is 1 February 2005.

If there are more candidates than open positions, ballots will be mailed to all members within 30 days after the close of nominations. Marked ballots must be postmarked no more than six (6) weeks following the mailing; the exact dates will be noted on the ballots.

Ballots will be counted by a disinterested party not part of the TUG organization. The results of the election should be available early in June, and will be announced in TUGboat as well as through various TEX-related electronic lists.

It is the intent of TUG that all wishing to participate in TUG elections do so. If you experience difficulty in submitting your nomination, do not receive your ballot in a timely way, or experience other problems, please contact the TUG office or the Elections Committee at election@tug.org.

Barbara Beeton

for the Elections Committee

Nomination Form

Only TUG members whose dues have been paid for 2005 will be eligible to participate in the election. The signatures of two (2) members in good standing at the time they sign the nomination form are required in addition to that of the nominee. Type or print names clearly; both supporting signatures are required. Please use the name by which you are known to TUG; all names must be identified from the TUG membership records before being accepted as valid. Please provide an email address if you wish an electronic confirmation of receipt of your nomination or supporting signature.

The undersigned TUG members propose the nomination of:

Print name of Nominee

signature Email or fax

date

for the position of (check one):

TUG President

Member of the TUG Board of Directors for a term beginning with the TUG 2005 Annual Meeting, August 2005.

Members supporting this nomination:

1.  _{please print name}_{signature}  _{Email or fax}_{date}

2.  _{please print name}_{signature}  _{Email or fax}_{date}

Return this nomination form to the TUG office (faxed forms will be accepted). Nomination forms and all required supplementary material (photograph, biography and personal statement for inclusion on the ballot) must be received in the TUG office no later than 1 February 2005.^{1} It is the responsibility of the candidate to ensure that this deadline is met; candidates are advised to check with the TUG office that their nomination forms are in order.

nomination form

biography/personal statement

photograph

TEX Users Group fax: +1 503 223-3960 Nominations for 2005 Election P.O. Box 2311 Portland, OR 97208–2311 USA

1Supplementary material may be sent separately from the form, and supporting signatures need not all appear on one form.